# A Tutte-type canonical decomposition of 3- and 4-connected graphs

### v2

@@ 2. Essential terminology
This section introduces all terminology used by Theorem˜1. For graph-theoretic terminology we follow [D24a].
@@ 2.1. Tetra-separations
The lower-bound $|G|\geqslant 8$ in Lemma˜2.3 (ii) is tight: the graph depicted in Figure˜2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.
@@ 2.2. Stars, mixed-tree-decompositions and torsos
The following definition is supported by Figure˜2.2. Let $G$ be a graph. A pair $\mathcal{T}=(T,(V_{t})_{t\in T})$ of a tree $T$ and a family of vertex sets $V_{t}\subseteq V(G)$ indexed by the nodes $t\in T$ is a mixed-tree-decomposition of $G$ if it satisfies the following two conditions:
@@
by making each of the vertex-sets $\lambda(\alpha_{\mathcal{T}}(s,t))$ into a clique. From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see Figure˜2.3.
@@
Figure 2.3. The torsos of the mixed-tree-decomposition depicted in Figure˜$2.$
@@ 3. Proofs of Theorem˜1 and ˜2 and strengthenings thereof
The outcome in Theorem˜3.1 is not in general unique. For example, consider the graph $G$ depicted in Figure˜3.1 on the left. Every tetra-separation of $G$ is crossed, hence $\sigma=\emptyset$ is a splitting star of the set $N(G)$ of all totally-nested tetra-separations of $G$. On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see Theorem˜3.1 ˜(ii). On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2. From their perspective, $G$ is a generalised wheel, see Theorem˜3.1 ˜(iii).
For another example, consider the graph $G$ depicted in Figure˜3.2 on the left. Again, $\sigma=\emptyset$ is a splitting star of $N(G)$. On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos, see Theorem˜3.1 ˜(ii). On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4. From their perspective, $G$ is a $K_{4,4}$, see Theorem˜3.1 ˜(iv).
@@ 4. How tetra-separations cross
We require the following terminology for corner diagrams from [CK23, §1.3], which is supported by Figure˜4.1. Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$. All of the following definitions depend on $(A,B)$ and $(C,D)$ being clear from context. The corner for $\{A,C\}$, or the corner $AC$ for short, is the vertex $(A\smallsetminus B)\cap(C\smallsetminus D)$. The corners for the other three combinations of one side from $(A,B)$ and one side from $(C,D)$ are defined analogously. Two corners are adjacent when their pair share a side, and opposite otherwise. An edge of $G$ is diagonal if its endvertices lie in opposite corners. The centre consists of all vertices in the vertex-centre and all diagonal edges.
@@
The following definition is supported by Figure˜4.2. The corner for $AC$ is potter if it is empty and its adjacent links have size two and there exists a path $u_{1}u_{2}u_{3}u_{4}$ in $G$ such that $u_{1}u_{2}$ dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and $u_{3}u_{4}$ dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link.
@@
A mixed-separation $(A,B)$ is an edge-cut if $S(A\cap B)\subseteq E(G)$. An edge-cut $(A,B)$ is atomic if $|A|=1$ or $|B|=1$.
@@ 6.3. Potter-linked
Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a graph $G$. Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see Figure˜4.2. Then $\{u_{2},u_{3}u_{4}\}$ is $A$-weird and $\{u_{3},u_{1}u_{2}\}$ is $C$-weird.
@@
[…]
The following definition is supported by Figure˜6.1.
@@
Note that this is not symmetric in $\pi_{1},\pi_{2}$. We say that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(A,B)$ if for both $X\in\{A,B\}$$\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $X$ and $\pi_{2}$ is 1-potter-linked to $\pi_{1}$ around $X$ with respect to $(A,B)$. A tetra-separation $(A,B)$ is 1-potter-linked if $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked for every balanced bipartition $\{\pi_{1},\pi_{2}\}$ of $S(A,B)$.
@@
By $\neg$ (6.9.3), there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$. By (Menger for independent paths). (6.2), there is a mixed-separation $(C_{1},D_{1})$ of $G[B]$ of order $\leqslant 2$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\cap B\subseteq C_{1}\smallsetminus D_{1}$ and $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}$. In particular, $|C_{1}\smallsetminus D_{1}|\geqslant 2$ and $|D_{1}\smallsetminus C_{1}|\geqslant 2$. See Figure˜6.2 for the next definition. We define $(C_{2},D_{2}):=(C_{1}\cup a_{1},(D_{1}\cup A)\smallsetminus\{v_{1}\})$. We have $S(C_{2},D_{2})=S(C_{1},D_{1})\cup\{a^{\prime}_{1}v_{1},a_{1}\}$ since $v_{1}$ does not have a neighbour in $A$ besides $a_{1}^{\prime}$. By 4-connectivity of $G$ and Lemma˜6.1, $(C_{2},D_{2})$ has order exactly 4. Let $(C,D)$ be the left-right-reduction of $(C_{2},D_{2})$. Since $C_{1}\smallsetminus D_{1}\subseteq C_{2}\smallsetminus D_{2}$ we have $|C_{2}\smallsetminus D_{2}|\geqslant 2$, and similarly $|D_{2}\smallsetminus C_{2}|\geqslant 2$. Hence $(C,D)$ is a tetra-separation by ˜5.10.
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 2-potter-linked with respect to $(A,B)$. Say (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$; that is to say, $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird. For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see Figure˜6.3. Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A\smallsetminus\{a_{1}\}$, and let $b_{2}^{\prime}$ be the unique neighbour of $v_{2}$ in $B\smallsetminus\{b_{1}\}$. By the degree-condition, we have $a^{\prime}_{1}\in A\smallsetminus B$ and $b^{\prime}_{1}\in B\smallsetminus A$.
@@
We combine Lemma˜6.3, Lemma˜6.5 and Corollary˜6.14 with the (Crossing Lemma). (4.1). ∎
@@ 7. Thickened $K_{4,m}$’s
By assumption, both $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ have two components. In particular, $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Then by Lemma˜5.4(A,B)(A,B) satisfies the matching-condition. Let $v$ be a vertex in $S(A,B)$. By 4-connectivity of $G$ and by Lemma˜6.1, $v$ has a neighbour in every component of GS(A,B)G-S(A,B). In particular, $v$ satisfies the degree-condition. Hence, $(A,B)$ is a tetra-separation. ∎
@@ 8.1. Block-bagel
The following setting is supported by Figure˜8.1.
@@ 8.2. N-Paths lemmata
The following two lemmata are supported by Figure˜8.2.
@@
The following lemma is supported by Figure˜8.3.
@@ 8.3. Totally-nested tetra-separations from good bags
The following setting is supported by Figure˜8.4.
@@
For each helpful $h\in V(O)\cup E(O)$ we construct a $z^{\prime}_{1}$$z^{\prime}_{2}$ path $P_{h}$, as follows. If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by Lemma˜8.7, and we let $P_{h}$ be the path defined by these two edges; see Figure˜8.5. If $h$ is a helpful node, then we use the (Bag-Path Lemma). (8.9) to let $P_{h}$ by a $z^{\prime}_{1}$$z^{\prime}_{2}$ path avoiding the edge $z^{\prime}_{1}z^{\prime}_{2}$ (if it exists) and with interior in $\mathring{G}_{h}$.
@@
The following lemma is supported by Figure˜8.6.
@@ 8.4. Splitting stars and torsos
[…]
The following two lemmata are supported by Figure˜8.7.
@@
If $d$ is a vertex, then we claim that every neighbour $u$ of $d$ in $A\smallsetminus B$ also also lies in $A^{\prime}\smallsetminus B^{\prime}$. Indeed, since $ud$ is not a dangling edge, $u$ is contained in the corner $AD$ or the $A$-link with respect to $(A,B),(C,D)$. By Lemma˜8.2 (i), no endvertex of $e$ lies in the $A$-link with respect to $(A,B),(C,D)$. So, we have $u\in A^{\prime}\smallsetminus B^{\prime}$.
@@ 9.1. Bananas
Let $G$ be a 4-connected graph. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross with all links of size two. Recall that the centre is empty by the (Crossing Lemma). (4.1). We obtain the graph $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{G^{*}}$} from $G$ by adding edges, as follows. For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see Figure˜9.1. We will only use $G^{*}$ when $(A,B)$ and $(C,D)$ are clear from context. The expanded corner for $AC$ is obtained from the corner $(A\smallsetminus B)\cap(C\smallsetminus D)$ for $AC$ by adding the endvertices of all edges contained in the two adjacent links. The banana for $AC$, or $AC$-banana, is the subgraph of $G^{*}$ induced by the expanded corner for $AC$. The expanded corners and bananas are defined similarly for the other three corners.
@@
$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii). Say $K$ is the corner for $AC$. Let $\beta$ denote the $AC$-banana. Let $e(A)$ denote the $AC$-edge for the $A$-link, and let $e(C)$ denote the $AC$-edge for the $C$-link. Assume that $e(A)=e(C)=:e$. We will show that $AC$ is potter. If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see Figure˜9.2. As $G$ is 4-connected, the corner for $AC$ cannot contain any other vertices besides the endvertices of $e$. But then both endvertices of $e$ have degree $\leqslant 3$ in $G$, contradicting that $G$ is 4-connected. So we may assume that $e$ has an endvertex $v$ that lies in the $A$-link, say. Then the $C$-link contains an edge $f$ that ends in $v$. Hence $f$ dangles from the $A$-link through the $C$-link, which implies that the corner for $AC$ is potter by the (Crossing Lemma). (4.1).

### v3

@@ 2. Essential terminology
This section introduces all terminology used by Theorem˜1. For graph-theoretic terminology we follow [R24].
@@ 2.1. Tetra-separations
The lower-bound $|G|\geqslant 8$ in lemma˜2.3 (ii) is tight: the graph depicted in fig.˜2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.
@@ 2.2. Stars, mixed-tree-decompositions and torsos
The following definition is supported by fig.˜2.2. Let $G$ be a graph. A pair $\mathcal{T}=(T,(V_{t})_{t\in T})$ of a tree $T$ and a family of vertex sets $V_{t}\subseteq V(G)$ indexed by the nodes $t\in T$ is a mixed-tree-decomposition of $G$ if it satisfies the following two conditions:
@@
by making each of the vertex-sets $\lambda(\alpha_{\mathcal{T}}(s,t))$ into a clique. From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see fig.˜2.3.
@@
Figure 2.3. The torsos of the mixed-tree-decomposition depicted in fig.˜$2.$
@@ 3. Proofs of Theorem˜1 and ˜2 and strengthenings thereof
The outcome in theorem˜3.1 is not in general unique. For example, consider the graph $G$ depicted in fig.˜3.1 on the left. Every tetra-separation of $G$ is crossed, hence $\sigma=\emptyset$ is a splitting star of the set $N(G)$ of all totally-nested tetra-separations of $G$. On the left of fig.˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see theorem˜3.1 ˜(ii). On the right of fig.˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2. From their perspective, $G$ is a generalised wheel, see theorem˜3.1 ˜(iii).
For another example, consider the graph $G$ depicted in fig.˜3.2 on the left. Again, $\sigma=\emptyset$ is a splitting star of $N(G)$. On the left of fig.˜3.2, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos, see theorem˜3.1 ˜(ii). On the right of fig.˜3.2, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4. From their perspective, $G$ is a $K_{4,4}$, see theorem˜3.1 ˜(iv).
@@ 4. How tetra-separations cross
We require the following terminology for corner diagrams from [CK23, §1.3], which is supported by fig.˜4.1. Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$. All of the following definitions depend on $(A,B)$ and $(C,D)$ being clear from context. The corner for $\{A,C\}$, or the corner $AC$ for short, is the vertex $(A\smallsetminus B)\cap(C\smallsetminus D)$. The corners for the other three combinations of one side from $(A,B)$ and one side from $(C,D)$ are defined analogously. Two corners are adjacent when their pair share a side, and opposite otherwise. An edge of $G$ is diagonal if its endvertices lie in opposite corners. The centre consists of all vertices in the vertex-centre and all diagonal edges.
@@
The following definition is supported by fig.˜4.2. The corner for $AC$ is potter if it is empty and its adjacent links have size two and there exists a path $u_{1}u_{2}u_{3}u_{4}$ in $G$ such that $u_{1}u_{2}$ dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and $u_{3}u_{4}$ dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link.
@@
A mixed-separation $(A,B)$ is an edge-cut if $S(A,B)\subseteq E(G)$. An edge-cut $(A,B)$ is atomic if $|A|=1$ or $|B|=1$.
@@ 6.3. Potter-linked
Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a graph $G$. Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see fig.˜4.2. Then $\{u_{2},u_{3}u_{4}\}$ is $A$-weird and $\{u_{3},u_{1}u_{2}\}$ is $C$-weird.
@@
We claim that $(C^{\prime},D^{\prime})$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links and no corner is potter. Since $\hat{\pi}_{1}\subseteq C\smallsetminus D\subseteq C^{\prime}\smallsetminus D^{\prime}$ and $\hat{\pi}_{2}\subseteq D\smallsetminus C\subseteq D^{\prime}\smallsetminus C^{\prime}$ by ˜5.9, $\pi_{1}$ forms the $C^{\prime}$-link while $\pi_{2}$ forms the $D^{\prime}$-link. In particular, $(C^{\prime},D^{\prime})$ crosses $(A,B)$. Assume for contradiction that some corner is potter, say the corner $AC^{\prime}$. Then $\pi_{1}$ is $A$-weird and hence contains an edge $e$. Let $a$ be the endvertex of $e$ in $A\smallsetminus B$. So, $a\in\hat{\pi}_{1}\subseteq C^{\prime}\smallsetminus D^{\prime}$. Therefore, $a$ is contained in the corner $AC^{\prime}$. In particular, the corner $AC^{\prime}$ is not empty, and so it is not potter, a contradiction. ∎
The following definition is supported by fig.˜6.1.
@@
Note that this is not symmetric in $\pi_{1},\pi_{2}$. We say that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(A,B)$ if for both $X\in\{A,B\}$, $\pi_{1}$ is 1-potter-linked to $\pi_{2}$ around $X$ and $\pi_{2}$ is 1-potter-linked to $\pi_{1}$ around $X$ with respect to $(A,B)$. A tetra-separation $(A,B)$ is 1-potter-linked if $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked for every balanced bipartition $\{\pi_{1},\pi_{2}\}$ of $S(A,B)$.
@@
By $\neg$ (6.9.3), there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$. By (Menger for independent paths). (6.2), there is a mixed-separation $(C_{1},D_{1})$ of $G[B]$ of order $\leqslant 2$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\cap B\subseteq C_{1}\smallsetminus D_{1}$ and $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}$. In particular, $|C_{1}\smallsetminus D_{1}|\geqslant 2$ and $|D_{1}\smallsetminus C_{1}|\geqslant 2$. See fig.˜6.2 for the next definition. We define $(C_{2},D_{2}):=(C_{1}\cup a_{1},(D_{1}\cup A)\smallsetminus\{v_{1}\})$. We have $S(C_{2},D_{2})=S(C_{1},D_{1})\cup\{a^{\prime}_{1}v_{1},a_{1}\}$ since $v_{1}$ does not have a neighbour in $A$ besides $a_{1}^{\prime}$. By 4-connectivity of $G$ and lemma˜6.1, $(C_{2},D_{2})$ has order exactly 4. Let $(C,D)$ be the left-right-reduction of $(C_{2},D_{2})$. Since $C_{1}\smallsetminus D_{1}\subseteq C_{2}\smallsetminus D_{2}$ we have $|C_{2}\smallsetminus D_{2}|\geqslant 2$, and similarly $|D_{2}\smallsetminus C_{2}|\geqslant 2$. Hence $(C,D)$ is a tetra-separation by ˜5.10.
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 2-potter-linked with respect to $(A,B)$. Say (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$; that is to say, $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird. For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see fig.˜6.3. Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A\smallsetminus\{a_{1}\}$, and let $b_{2}^{\prime}$ be the unique neighbour of $v_{2}$ in $B\smallsetminus\{b_{1}\}$. By the degree-condition, we have $a^{\prime}_{1}\in A\smallsetminus B$ and $b^{\prime}_{1}\in B\smallsetminus A$.
@@ 7. Thickened $K_{4,m}$’s
By assumption, both $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ have two components. In particular, $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Then by lemma˜5.4, $(A,B)$ satisfies the matching-condition. Let $v$ be a vertex in $S(A,B)$. By 4-connectivity of $G$ and by lemma˜6.1, $v$ has a neighbour in every component of GS(A,B)G-S(A,B). In particular, $v$ satisfies the degree-condition. Hence, $(A,B)$ is a tetra-separation. ∎
@@ 8.1. Block-bagel
The following setting is supported by fig.˜8.1.
@@ 8.2. N-Paths lemmata
The following two lemmata are supported by fig.˜8.2.
@@
The following lemma is supported by fig.˜8.3.
@@ 8.3. Totally-nested tetra-separations from good bags
The following setting is supported by fig.˜8.4.
@@
For each helpful $h\in V(O)\cup E(O)$ we construct a $z^{\prime}_{1}$$z^{\prime}_{2}$ path $P_{h}$, as follows. If $h$ is an edge, then its adhesion-vertex send edges to both $z^{\prime}_{i}$ by lemma˜8.7, and we let $P_{h}$ be the path defined by these two edges; see fig.˜8.5. If $h$ is a helpful node, then we use the (Bag-Path Lemma). (8.9) to let $P_{h}$ by a $z^{\prime}_{1}$$z^{\prime}_{2}$ path avoiding the edge $z^{\prime}_{1}z^{\prime}_{2}$ (if it exists) and with interior in $\mathring{G}_{h}$.
@@
The following lemma is supported by fig.˜8.6.
@@ 8.4. Splitting stars and torsos
By ˜8.27, $\sigma(\mathcal{O})$ is a set of totally-nested tetra-separations. It remains to show that for all good nodes $s\neq t$, we have $(U_{s},W_{s})\leqslant(W_{t},U_{t})$. We have $U_{s}\subseteq V(G_{s})$ and $U_{t}\subseteq V(G_{t})$. Moreover, by ˜8.15, for all nodes $r\in V(O-s)$ that have a 2-connected bag, we have $V(G_{r})\subseteq W_{s}$. Therefore, $U_{t}\subseteq V(G_{t})\subseteq W_{s}$. Similarly, $U_{s}\subseteq V(G_{s})\subseteq W_{t}$. It follows that $(U_{s},W_{s})\leqslant(W_{t},U_{t})$. ∎
The following two lemmata are supported by fig.˜8.7.
@@
If $d$ is a vertex, then we claim that every neighbour $u$ of $d$ in $A\smallsetminus B$ also lies in $A^{\prime}\smallsetminus B^{\prime}$. Indeed, since $ud$ is not a dangling edge, $u$ is contained in the corner $AD$ or the $A$-link with respect to $(A,B),(C,D)$. By lemma˜8.2 (i), no endvertex of $e$ lies in the $A$-link with respect to $(A,B),(C,D)$. So, we have $u\in A^{\prime}\smallsetminus B^{\prime}$.
@@ 9.1. Bananas
Let $G$ be a 4-connected graph. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross with all links of size two. Recall that the centre is empty by the (Crossing Lemma). (4.1). We obtain the graph $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{G^{*}}$} from $G$ by adding edges, as follows. For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see fig.˜9.1. We will only use $G^{*}$ when $(A,B)$ and $(C,D)$ are clear from context. The expanded corner for $AC$ is obtained from the corner $(A\smallsetminus B)\cap(C\smallsetminus D)$ for $AC$ by adding the endvertices of all edges contained in the two adjacent links. The banana for $AC$, or $AC$-banana, is the subgraph of $G^{*}$ induced by the expanded corner for $AC$. The expanded corners and bananas are defined similarly for the other three corners.
@@
$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii). Say $K$ is the corner for $AC$. Let $\beta$ denote the $AC$-banana. Let $e(A)$ denote the $AC$-edge for the $A$-link, and let $e(C)$ denote the $AC$-edge for the $C$-link. Assume that $e(A)=e(C)=:e$. We will show that $AC$ is potter. If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see fig.˜9.2. As $G$ is 4-connected, the corner for $AC$ cannot contain any other vertices besides the endvertices of $e$. But then both endvertices of $e$ have degree $\leqslant 3$ in $G$, contradicting that $G$ is 4-connected. So we may assume that $e$ has an endvertex $v$ that lies in the $A$-link, say. Then the $C$-link contains an edge $f$ that ends in $v$. Hence $f$ dangles from the $A$-link through the $C$-link, which implies that the corner for $AC$ is potter by the (Crossing Lemma). (4.1).